# A Neuromorphic Architecture for Scalable Event-Based Control

### v1

@@ 9 Conclusion
There are many questions that require research beyond the material of this paper. Primarily and foremost, we wish to develop learning and adaptation algorithms that can take advantage of the proposed neuromorphic architecture. The event-based nature of the neural network suggests the potential of Hebbian rules for learning and adaptation. Another key question is the hardware realisation of the ”nervous system” architecture., and the respective advantages of digital and analog devices all the way from sensing to actuation. Those questions will be addressed in a future publication.
@@ 10.1 Parameters for simulations and implementation detail for snake controller
All spiking neurons used in this paper has the following parameters: $\tau_{s}=20$, $\alpha_{f}^{-}=\alpha_{s}^{+}=2$, $R=0.5$, ${\bar{V}}=-1.5$. For bursting neuron in Figure 4, we use: $\tau_{s}=5$, $\tau_{us}=100,\alpha_{f}^{-}=\alpha_{s}^{+}=2$, $\alpha_{s}^{-}=1.5$, $R=0.5,{\bar{V}}_{f}=3$, ${\bar{V}}_{s}=-1.5$. We use $\alpha_{us}^{+}=1.5$ for trajectory A and $\alpha_{us}^{+}=1.7$ for trajectory B.

### v2

@@ 9 Conclusion
There are many questions that require research beyond the material of this paper. Primarily and foremost, we wish to develop learning and adaptation algorithms that can take advantage of the proposed neuromorphic architecture. The event-based nature of the neural network suggests the potential of Hebbian rules for learning and adaptation. Another key question is the hardware realisation of the ”nervous system” architecture, and the respective advantages of digital and analog devices all the way from sensing to actuation. Those questions will be addressed in a future publication.
@@ 10.1 Parameters for simulations and implementation detail for snake controller
All spiking neurons used in this paper has the following parameters: $\tau_{s}=20$, $\alpha_{f}^{-}=\alpha_{s}^{+}=2$, $R=2$, ${\bar{V}}=-1.5$. For bursting neuron in Figure 4, we use: $\tau_{s}=5$, $\tau_{us}=100,\alpha_{f}^{-}=\alpha_{s}^{+}=2$, $\alpha_{s}^{-}=1.5$, $R=0.5,{\bar{V}}_{f}=3$, ${\bar{V}}_{s}=-1.5$. We use $\alpha_{us}^{+}=1.5$ for trajectory A and $\alpha_{us}^{+}=1.7$ for trajectory B.